OpenStudy (anonymous):

metal will be givin What is the slope of the line that passes through the points (–2, 5) and (1, 11)? a.-2 b.1/2 c.2 d.-6

OpenStudy (the_fizicx99):

M = y2-y1/x2-x1 M = 11-5/1-2 6/-3 Solve that na dthat'll give you your M or Slope

OpenStudy (the_fizicx99):

and* that'll*

OpenStudy (marigirl):

slope refers to the gradient which can be obtained by using the formula\[\frac{ y _{2}-y _{1} }{ x _{2}-x{1} }\]